Find the regression equation, letting the first variable be the predictor x) variable. Using the listed actress/actor ages inPart b) Th best predicted age of the best actor winner given that the age of the best actress winner is 29 years is ___ years old.
Answer #1
X Y XY X^2 Y^2
27 44 1188 729 1936
29 38 1102 841 1444
29 37 1073 841 1369
59 44 2596 3481 1936
30 51 1530 900 2601
31 46 1426 961 2116
43 62 2666 1849 3844
29 52 1508 841 2704
64 37 2368 4096 1369
22 54 1188 484 2916
43 44 1892 1849 1936
55 34 1870 3025 1156

PREDICTED Y= a+b X . SDy a = Y - bX

From the above table and formula we get the value are as:

n 12
sum(XY) 20407.00
sum(X) 461.00
sum(Y) 543.00
sum(X^2) 19897.00
sum(Y^2) 25327.00
b -0.207
a 53.2

ycap = a +bx

ycap = 53.2 - 0.207x

when x = 29

Predicted value = 53.2 - 0.207 * 29

= 47.20

= approx 47 years old